# No Compromise in Solution Quality: Speeding Up Belief-dependent Continuous POMDPs via Adaptive Multilevel Simplification

### v1

@@ Abstract
Continuous POMDPs with general belief-dependent rewards are notoriously difficult to solve online. In this paper, we present a complete provable theory of adaptive multilevel simplification for the setting of a given externally constructed belief tree and MCTS that constructs the belief tree on the fly using an exploration technique. Our theory allows to accelerate POMDP planning with belief-dependent rewards without any sacrifice in the quality of the obtained solution. We rigorously prove each theoretical claim in the proposed unified theory. Using the general theoretical results, we present three algorithms to accelerate continuous POMDP online planning with belief-dependent rewards. Our two algorithms, SITH-BSP and LAZY-SITH-BSP, can be utilized on top of any method that constructs a belief tree externally. The third algorithm, SITH-PFT, is an anytime MCTS method that permits to plug-in any exploration technique. All our methods are guaranteed to return exactly the same optimal action as their unsimplified equivalents. We replace the costly computation of information-theoretic rewards with novel adaptive upper and lower bounds which we derive in this paper, and are of independent interest. We show that they are easy to calculate and can be tightened by the demand of our algorithms. Our approach is general; namely, any bounds that monotonically converge to the reward can be easily plugged-in to achieve significant speedup without any loss in performance. Our theory and algorithms support the challenging setting of continuous states, actions, and observations. The beliefs can be parametric or general and represented by weighted particles. We demonstrate in simulation a significant speedup in planning compared to baseline approaches with guaranteed identical performance.
@@ 1 Introduction
Efficiently solving Partially Observable Markov Decision Processes (POMDPs) implies enabling autonomous agents and robots to plan under uncertainty (Smith and Simmons, 2004; Kurniawati et al., 2008; Silver and Veness, 2010; Ye et al., 2017; Sunberg and Kochenderfer, 2018; Garg et al., 2019). Typical sources of uncertainty are the imprecise actions, sensor type, sensor noise, imprecise models, and unknown agent surroundings. However, solving a POMDP is notoriously hard. Specifically, it was proven to be PSPACE-complete (Papadimitriou and Tsitsiklis, 1987).
The actual POMDP state is hidden. Instead, at each time step, the robot shall decide which action to take based on the distribution over the state, given the corresponding history of performed actions and observations received so far. Such a distributions received the name “beliefs”. In a planning session, the robot has to take into account all possible future actions interleaved with possible observations. Each such future history of the length of predefined horizon defines a lace of the future beliefs (blue lace in Fig. 1) and corresponding cumulative rewards named return. Solving POMDP in the most common sense means finding a mapping from belief to action called policy, which maximizes expected return.
Earlier offline solvers such as (Smith and Simmons, 2004; Kurniawati et al., 2008) are applicable to small or moderately sized discrete POMDP. These methods require passage over all possible states and observations (Kochenderfer et al., 2022) since they are built on value iteration of $\alpha$-vectors, so called full-width methods (Silver and Veness, 2010). More recent online solvers are suitable for POMDPs with large but discrete action, state, and observation spaces (Ye et al., 2017; Silver and Veness, 2010). Still, continuous state, action, and observation spaces remain to be an open problem (Sunberg and Kochenderfer, 2018). Another challenging aspect of solving POMDP and the subject of interest in this paper is general belief distributions represented by weighted particles. Further in the manuscript we will regard the combination of both, nonparametric beliefs and fully continuous POMDP as nonparametric fully continuous setting.
@@
The reward function in a classical POMDP is assumed to have a specific structure, namely, to be the expectation with respect to the belief of the state-dependent reward function. While alleviating the solution, this formulation does not support more general, belief-dependent reward functions, such as information-theoretic rewards. However, POMDP planning with belief-dependent rewards is essential for various problems in robotics and Artificial Intelligence (AI), such as informative planning (Hollinger and Sukhatme, 2014), active localization (Burgard et al., 1997), active Simultaneous Localization and Mapping (SLAM) (Stachniss et al., 2005), Belief Space Planning (BSP) (Indelman et al., 2015; Van Den Berg et al., 2012; Platt et al., 2010).
@@
Let us focus for the moment on differential entropy. Even if the belief is parametric but not Gaussian, calculating the exact value of differential entropy involves intractable integrals. This fact also motivates to use a weighted particles representation for the belief. In this case differential entropy can be estimated, for instance by Kernel Density Estimation (KDE) (Fischer and Tas, 2020) or a model-based estimator (Boers et al., 2010). However, these estimators have quadratic cost in the number of samples and are usually the bottleneck of planning algorithms. The reason is that this increased computational burden is incurred for all nodes in the belief tree. Importantly, the estimation errors of these estimators with respect to differential entropy over theoretical belief are out of the reach due to the unavailability of both, the theoretical belief and the entropy on top of it. Yet, common estimators are converging due to the convergence (almost surely (Crisan and Doucet, 2002)) of the belief represented by particles to the theoretical belief. This prompts us to use as many belief particles as possible to get closer to the theoretical belief. Increasing number of belief particles severely impacts planning time.
@@ 1.1 Related Work
Allowing general belief-dependent rewards in POMDP while solving such a problem efficiently is a long standing effort. Some previous seminal works such as $\rho$-POMDP (Araya et al., 2010; Fehr et al., 2018) as well as (Dressel and Kochenderfer, 2017) have focused on discrete domains, small sized spaces and have tackled the offline solvers. Furthermore, these approaches are limited to piecewise linear and convex or Lipschitz-continuous rewards. Another work named POMDP-IR (Spaan et al., 2015) suggest an interesting framework for specific form of information rewards involving manipulations on the action space. Still, in (Araya et al., 2010; Fehr et al., 2018; Dressel and Kochenderfer, 2017) the state, action and observation spaces are discrete and small sized. Another line of works is Belief Space Planning (BSP) (Platt et al., 2010; Van Den Berg et al., 2012; Indelman et al., 2015). These approaches are designed for fully continuous POMDPS, but limited to Gaussian beliefs. In striking contrast, our approach is centered in the more challenging fully continuous domain and nonparametric general beliefs represented by particles.
One way to tackle a nonparamteric fully continuous setting with belief dependent reward is to reformulate POMDP as a Belief-MDP (BMDP). On top of this reformulation one can utilize MDP sampling based methods such as Sparse Sampling (SS) proposed by Kearns et al. (2002). However, this algorithm still suffers from the curse of history and such that increasing the horizon is still problematic.
@@
MCTS algorithms assume that calculating the reward over the belief node does not pose any computational difficulty. Information-theoretic rewards violate this assumption. When the reward is a general function of the belief, the origin of the computational burden is shifted towards the reward calculation. Moreover, belief-dependent rewards require the complete set of belief particles at each node in the belief tree. Therefore, algorithms such as POMCP (Silver and Veness, 2010), and its numerous predecessors are inapplicable since they simulate each time a single particle down the tree when expanding it. DESPOT based algorithms behave similarly (Ye et al., 2017), with the DESPOT-$\alpha$ as an exception (Garg et al., 2019). DESPOT-$\alpha$ simulates a complete set of particles. However, the DESPOT-$\alpha$ tree is built using $\alpha$-vectors, such that they are an indispensable part of the algorithm. The standard $\alpha$-vectors technique requires that the reward is state dependent, and the reward over the belief is merely expectation over the state reward. In other words DESPOT-$\alpha$ does not support belief-dependent rewards since it contradicts the application of the $\alpha$-vectors.
@@
The only approach posing no restrictions on the structure of belief-dependent reward and not suffering from limiting assumptions is Particle Filter Tree (PFT). The idea behind PFT is to apply MCTS over Belief-MDP (BMDP). The authors of (Sunberg and Kochenderfer, 2018) augmented PFT with Double Progressive Widening (DPW) to support continuous spaces in terms of actions, states and observations, and coined the name PFT-DPW. PFT-DPW utilizes the UCB strategy and maintains a complete belief particle set at each belief tree node. Recently, Fischer and Tas (2020) presented Information Particle Filter Tree (IPFT), a method to incorporate information-theoretic rewards into PFT. The IPFT simulates small subsets of particles sampled from the root of the belief tree and averages entropies calculated over these subsets, enjoying a fast runtime. However, differential entropy estimated from a small-sized particle set can be significantly biased. This bias is unpredictable and unbounded, therefore, severely impairs the performance of the algorithm. In other words, celerity comes at the expense of quality. Oftentimes, the policy defined by this algorithm is very far from optimal given a time budget. (Fischer and Tas, 2020) provides guarantees solely for the asymptotic case, i.e, the number of state samples (particles) tends to infinity. Asymptotically their algorithm behaves precisely as the PFT-DPW in terms of running speed and performance. Yet, in practice the performance of IPFT in terms of optimality can degrade severely compared to PFT-DPW. Moreover, (Fischer and Tas, 2020) does not provide any study of comparison of IPFT against PFT-DPW with an information-theoretic reward. Prompted by this insight, we chose the PFT-DPW as our baseline approach, which we aim to accelerate. In contrast to IPFT designed specifically for differential entropy, our approach is suitable for any belief dependent reward and explicitly guarantees an identical solution to PFT-DPW with an information-theoretic reward, for any size of particle set representing the belief and serving as input to PFT-DPW.
@@
Recently, the novel paradigm of simplification has appeared in literature (Zhitnikov and Indelman, 2022; Sztyglic and Indelman, 2022; Elimelech and Indelman, 2022; Shienman and Indelman, 2022). The simplification is concerned with carefully replacing the nonessential elements of the decision making problem and quantifying the impact of this relaxation. Specifically, simplification methods are accompanied by stringent guarantees. A prominent aspect of a simplification paradigm is the usage of the bounds over the reward or the objective function. As opposed to approximations, the simplification framework always keeps some sort of connection to the original unsimplified problem and by that provides deterministic guarantees relative to the given solver. In spite that various objective function bounds have been practiced in (Ye et al., 2017; Smith and Simmons, 2004; Walsh et al., 2010; Kochenderfer et al., 2022), these techniques are not applicable in the realm of belief-dependent rewards and a fully continuous setting. In addition commonly these approaches assume that state dependent reward is trivially bounded by some constant.
@@ 1.2 Contributions
This work is about accelerating online decision making while obtaining exactly the same solution as without acceleration. Specifically, we contribute an adaptive multi-level simplification framework considering belief-dependent rewards, possibly nonparametric beliefs, and continuous state, observation and action spaces. We further develop particular simplification adaptation mechanisms for two main settings: the belief tree is given, or its construction is not coupled with the solution (e.g. SS), and an anytime setting (MCTS) where the belief tree construction is coupled with the solution due to an exploration strategy (e.g. UCB). In both cases, we guarantee no sacrifice in planning performance with a significant planning speedup.
This paper is an extension of the work presented in (Sztyglic and Indelman, 2022), which proposed novel adaptive bounds on the differential entropy estimator of (Boers et al., 2010) and introduced the simplification paradigm in the context of a given belief tree. For clarity, we list down the contributions of this work, in the order they are presented in the manuscript.
@@
    We derive novel lightweight adaptive bounds on the differential entropy estimator of (Boers et al., 2010) and prove the bounds presented are monotonic and convergent. Moreover, these bounds can be incrementally tightened. We believe these bounds are of interest on their own. The bounds are calculated using the simplified belief (See Fig. 1)

### v2

@@ Abstract
Continuous POMDPs with general belief-dependent rewards are notoriously difficult to solve online. In this paper, we present a complete provable theory of adaptive multilevel simplification for the setting of a given externally constructed belief tree and MCTS that constructs the belief tree on the fly using an exploration technique. Our theory allows to accelerate POMDP planning with belief-dependent rewards without any sacrifice in the quality of the obtained solution. We rigorously prove each theoretical claim in the proposed unified theory. Using the general theoretical results, we present three algorithms to accelerate continuous POMDP online planning with belief-dependent rewards. Our two algorithms, SITH-BSP and LAZY-SITH-BSP, can be utilized on top of any method that constructs a belief tree externally. The third algorithm, SITH-PFT, is an anytime MCTS method that permits to plug-in any exploration technique. All our methods are guaranteed to return exactly the same optimal action as their unsimplified equivalents. We replace the costly computation of information-theoretic rewards with novel adaptive upper and lower bounds which we derive in this paper, and are of independent interest. We show that they are easy to calculate and can be tightened by the demand of our algorithms. Our approach is general; namely, any bounds that monotonically converge to the reward can be utilized to achieve significant speedup without any loss in performance. Our theory and algorithms support the challenging setting of continuous states, actions, and observations. The beliefs can be parametric or general and represented by weighted particles. We demonstrate in simulation a significant speedup in planning compared to baseline approaches with guaranteed identical performance.
@@ 1 Introduction
[…]
[findent=2pt]Efficiently solving Partially Observable Markov Decision Processes (POMDPs) implies enabling autonomous agents and robots to plan under uncertainty (Smith and Simmons, 2004; Kurniawati et al., 2008; Silver and Veness, 2010; Ye et al., 2017; Sunberg and Kochenderfer, 2018; Garg et al., 2019). Typical sources of uncertainty are the imprecise actions, sensor type, sensor noise, imprecise models, and unknown agent surroundings. However, solving a POMDP is notoriously hard. Specifically, it was proven to be PSPACE-complete (Papadimitriou and Tsitsiklis, 1987).
The actual POMDP state is hidden. Instead, at each time step, the robot shall decide which action to take based on the distribution over the state, given the corresponding history of performed actions and observations received so far. Such a distribution received the name “belief”. In a planning session, the robot has to take into account all possible future actions interleaved with possible observations. Each such future history of the length of predefined horizon defines a lace of the future beliefs (blue lace in Fig. 1) and corresponding cumulative rewards named return. Solving POMDP in the most common sense means finding a mapping from belief to action called policy, which maximizes the expected return.
Earlier offline solvers such as (Smith and Simmons, 2004; Kurniawati et al., 2008) are applicable to small or moderately sized discrete POMDP. These methods require passage over all possible states and observations (Kochenderfer et al., 2022) since they are built on value iteration of $\alpha$-vectors, so called full-width methods (Silver and Veness, 2010). More recent online solvers are suitable for POMDPs with large but discrete action, state, and observation spaces (Ye et al., 2017; Silver and Veness, 2010). Still, continuous state, action, and observation spaces remain to be an open problem (Sunberg and Kochenderfer, 2018). Another challenging aspect of solving POMDP and the subject of interest in this paper is general belief distributions represented by weighted particles. Further in the manuscript we will regard the combination of both, nonparametric beliefs and a fully continuous POMDP as a nonparametric fully continuous setting.
@@
The reward function in a classical POMDP is assumed to have a specific structure, namely, to be the expectation with respect to the belief of the state-dependent reward function. While alleviating the solution, this formulation does not support more general, belief-dependent reward functions, such as information-theoretic rewards.
However, POMDP planning with belief-dependent rewards is essential for various problems in robotics and Artificial Intelligence (AI), such as informative planning (Hollinger and Sukhatme, 2014), active localization (Burgard et al., 1997), active Simultaneous Localization and Mapping (SLAM) (Stachniss et al., 2005), Belief Space Planning (BSP) (Indelman et al., 2015; Van Den Berg et al., 2012; Platt et al., 2010). The authors of (Araya et al., 2010) provide an extensive motivation for general belief-dependent rewards. One of the widely used such rewards is Information Gain, which involves the difference between differential entropies of two consecutive in time beliefs. Such a reward is crucial in exploration tasks because, in these tasks, the robot’s goal is to decrease uncertainty over the belief. For instance, uncertainty measures such as differential entropy and determinant of the covariance matrix of the belief cannot be represented as expectation over a state-dependent reward with respect to the belief. Another example of a belief-dependent reward is entropy over discrete variables that correspond to data association hypotheses (Pathak et al., 2018). Computationally-efficient information-theoretic BSP approaches have been investigated in recent years, considering Gaussian distributions (Kopitkov and Indelman, 2017, 2019; Elimelech and Indelman, 2022; Kitanov and Indelman, 2024).
@@
Let us focus for the moment on differential entropy. Even if the belief is parametric but not Gaussian, calculating the exact value of differential entropy involves intractable integrals. This fact also motivates to use a weighted particles representation for the belief. In this case differential entropy can be estimated, for instance by Kernel Density Estimation (KDE) (Fischer and Tas, 2020) or a model-based estimator (Boers et al., 2010). However, these estimators have quadratic cost in the number of samples and are usually the bottleneck of planning algorithms. The reason is that this increased computational burden is incurred for all nodes in the belief tree. Importantly, the estimation errors of these estimators with respect to differential entropy over theoretical belief are out of the reach due to the unavailability of both, the theoretical belief and the entropy on top of it. Yet, due to the convergence of the belief represented by particles to the theoretical belief (almost sure convergence (Crisan and Doucet, 2002)), the mentioned above estimators converge to the exact differential entropy. This prompts us to use as many belief particles as possible to get closer to the theoretical belief. Nevertheless, increasing the number of belief particles severely impacts planning time.
@@ 1.1 Related Work
Allowing general belief-dependent rewards in POMDP while solving such a problem efficiently is a long standing effort. Some previous seminal works such as $\rho$-POMDP (Araya et al., 2010; Fehr et al., 2018) as well as (Dressel and Kochenderfer, 2017) have focused on discrete domains, small sized spaces and have tackled the offline solvers. Furthermore, these approaches are limited to piecewise linear and convex or Lipschitz-continuous rewards. Another work named POMDP-IR (Spaan et al., 2015) suggest an interesting framework for specific form of information rewards involving manipulations on the action space. Still, in (Araya et al., 2010; Fehr et al., 2018; Dressel and Kochenderfer, 2017) the state, action and observation spaces are discrete and small sized. Another line of works is Belief Space Planning (BSP) (Platt et al., 2010; Van Den Berg et al., 2012; Indelman et al., 2015). These approaches are designed for fully continuous POMDPs, but limited to Gaussian beliefs. In striking contrast, our approach is centered in the more challenging fully continuous domain and nonparametric general beliefs represented by particles while at the same time our framework is general and supports also exact parametric beliefs.
One way to tackle a nonparametric fully continuous setting with belief dependent rewards is to reformulate POMDP as a Belief-MDP (BMDP). On top of this reformulation one can utilize MDP sampling based methods such as Sparse Sampling (SS) proposed by Kearns et al. (2002). However, this algorithm still suffers from the curse of history and such that increasing the horizon is still problematic.
@@
MCTS algorithms assume that calculating the reward over the belief node does not pose any computational difficulty. Information-theoretic rewards violate this assumption. When the reward is a general function of the belief, the origin of the computational burden is shifted towards the reward calculation. Moreover, in a non-parametric setting, belief-dependent rewards require a complete set of belief particles at each node in the belief tree. Therefore, algorithms such as POMCP (Silver and Veness, 2010), and its numerous predecessors are inapplicable since they simulate each time a single particle down the tree when expanding it. DESPOT based algorithms behave similarly (Ye et al., 2017), with the DESPOT-$\alpha$ as an exception (Garg et al., 2019). DESPOT-$\alpha$ simulates a complete set of particles. However, the DESPOT-$\alpha$ tree is built using $\alpha$-vectors, such that they are an indispensable part of the algorithm. The standard $\alpha$-vectors technique requires that the reward is state dependent, and the reward over the belief is merely expectation over the state reward. In other words DESPOT-$\alpha$ does not support belief-dependent rewards since it contradicts the application of the $\alpha$-vectors.
@@
The only approach posing no restrictions on the structure of belief-dependent reward and not suffering from limiting assumptions is Particle Filter Tree (PFT). The idea behind PFT is to apply MCTS over Belief-MDP (BMDP). The authors of (Sunberg and Kochenderfer, 2018) augmented PFT with Double Progressive Widening (DPW) to support continuous spaces in terms of actions, states and observations, and coined the name PFT-DPW. PFT-DPW utilizes the UCB strategy and maintains a complete belief particle set at each belief tree node. Recently, Fischer and Tas (2020) presented Information Particle Filter Tree (IPFT), a method to incorporate information-theoretic rewards into PFT. The IPFT simulates small subsets of particles sampled from the root of the belief tree and averages entropies calculated over these subsets, enjoying a fast runtime. However, differential entropy estimated from a small-sized particle set can be significantly biased. This bias is unpredictable and unbounded, therefore, severely impairs the performance of the algorithm. In other words, celerity comes at the expense of quality. Oftentimes, the policy defined by this algorithm is very far from optimal given a time budget. Fischer and Tas (2020) provides guarantees solely for the asymptotic case, i.e, the number of subsampled from the root belief state samples (particles) tends to infinity. Asymptotically their algorithm behaves precisely as the PFT-DPW in terms of running speed and performance. Yet, in practice the performance of IPFT in terms of optimality can degrade severely compared to PFT-DPW. Moreover, Fischer and Tas (2020) does not provide any study of comparison of IPFT against PFT-DPW with an information-theoretic reward. Another undesired characteristic of IPFT is that the averaging of the differential entropies is done implicitly and the number of averaged entropies per belief is the visitation count of the corresponding belief. Therefore, to properly compare IPFT with PFT-DPW one shall increase the number of simulations inside IPFT algorithm. We explain this aspect more thoroughly in Section 8.3.5. Prompted by these insights, we chose the PFT-DPW as our baseline approach, which we aim to accelerate. In contrast to IPFT designed specifically for differential entropy, our approach is suitable for any belief dependent reward and explicitly guarantees an identical solution to PFT-DPW with an information-theoretic reward, for any size of particle set representing the belief and serving as input to PFT-DPW.
@@
Recently, the novel paradigm of simplification has appeared in literature (Zhitnikov and Indelman, 2022b; Barenboim and Indelman, 2022, 2023; Zhitnikov and Indelman, 2024; Sztyglic and Indelman, 2022; Elimelech and Indelman, 2022; Shienman and Indelman, 2022; Kitanov and Indelman, 2024; Lev-Yehudi et al., 2024). The simplification is concerned with carefully replacing the nonessential elements of the decision making problem and quantifying the impact of this relaxation. Specifically, simplification methods are accompanied by stringent guarantees. A prominent aspect of a simplification paradigm is the usage of the bounds over the reward or the objective function. As opposed to approximations, the simplification framework always keeps some sort of connection to the original unsimplified problem and by that provides deterministic guarantees relative to the given solver. Despite that various objective function bounds have been practiced in (Ye et al., 2017; Smith and Simmons, 2004; Walsh et al., 2010; Kochenderfer et al., 2022), these techniques are not applicable in the realm of belief-dependent rewards and a fully continuous setting. In addition commonly these approaches assume that the state dependent reward is trivially bounded from below and above by some constant.
@@ 1.2 Contributions
This work is about accelerating online decision making while obtaining exactly the same solution as without acceleration. Specifically, we contribute an adaptive multi-level simplification framework that accounts for belief-dependent rewards, possibly nonparametric beliefs, and continuous state, observation and action spaces.
Our framework accepts as input adaptive monotonical and computationally inexpensive bounds over the exact or estimated reward. Given such reward bounds, it accelerates online decision-making. Specifically, given such adaptive monotonical reward bounds, it is possible to adaptively bound the value function for any given policy and expedite decision-making. If the value function bounds for different candidate policies do not overlap, we do not pay in terms of quality, namely, we obtain the same solution as the equivalent unsimplified method. In the case these bounds do overlap, then we can progressively tighten them by invoking a process that we shall call simplification adaptation or resimplification until they no longer overlap.
Our techniques return exactly the same solution as the unsimplified equivalent. Such an unsimplified baseline can correspond to decision-making problems where the reward can be exactly calculated (analytically), or where the reward is estimated. In either case, if the bounds over the corresponding reward are provided and satisfy the assumptions stated in Section 3.3, one can apply our framework to speedup the decision making process while obtaining the same best action as with the original rewards instead of the bounds. Such a capability is therefore particularly appealing in light of the information-theoretic rewards that are essential in numerous problems in robotics, but are often the computational bottleneck.
Further, there are two settings that we separately and explicitly discuss in this paper. We start from a given belief tree, that can be constructed by a POMDP solver that is not coupled with the solution, e.g., SS. In this setting we can prune branches of the belief tree whenever the mentioned objective bounds for different candidate policies or actions do not overlap.
We then discuss an anytime setting of MCTS, where the belief tree construction is coupled with the solution due to an exploration strategy (e.g. UCB). The exploration strategy builds upon an exploratory objective. Since the exploratory objective typically requires access to the objective estimates to select an action at each arrival to a belief node, we cannot prune suboptimal candidate actions. Instead, we can only dismiss them until the next arrival to this belief node. The simplification and reward bounds are used here to bound the exploratory objective and the value function at the root of the belief tree.
Finally, we focus on a specific simplification of nonparametric beliefs represented by particles and a differential entropy estimator as the reward function. Our simplification is subsampling of the original belief to a smaller sample size. We contribute novel computationally cheaper bounds over the differential entropy estimator on top of such a simplified belief and incorporate these bounds into our framework. By that we produce a specific embodiment of the general framework presented earlier.
To summarize, we list down the contributions of this work, in the order they are presented in the manuscript.
@@
    We derive novel lightweight adaptive bounds on the differential entropy estimator of (Boers et al., 2010) and prove the bounds presented are monotonic and convergent. Moreover, these bounds can be incrementally tightened. We believe these bounds are of interest on their own. The bounds are calculated using the simplified belief (See Fig. 1). We emphasize that any other bounds fulfilling assumptions declared in Section 3.3 can be utilized within our framework.
@@
This paper is an extension of the work presented in (Sztyglic and Indelman, 2022), which proposed novel adaptive bounds on the differential entropy estimator of (Boers et al., 2010) and introduced the simplification paradigm in the context of a given belief tree. To be precise we explicitly clarify how this work differs from the conference version of this paper (Sztyglic and Indelman, 2022). In this version, we extend the simplification framework to the rewards depending on a pair of consecutive-in-time beliefs, e.g., Information Gain as opposed to the conference version where such an extension was only mentioned. In this version, we provide alternative proof of these bounds and prove that these reward bounds are monotonic. In the setting of a given belief tree we present an additional algorithm, that we call LAZY-BSP. This algorithm is faster than SITH-BSP suggested in (Sztyglic and Indelman, 2022). Importantly, we extend our simplification framework to support also anytime MCTS planners. Additionally, we provide extensive performance evaluation of our methods in simulations.